# Cosmological Roles of Dark Photons in Axion-induced Electroweak Baryogenesis

### v2

@@ II.2 Constraints on axion
This contribution remains below the ACME II and JILA limits Andreev et al. (2018); Roussy et al. (2023) for $f_{\phi}$ larger than about $10$ TeV. Notably, many other electroweak baryogenesis scenarios face stringent EDM constraints, as achieving a strong first-order phase transition typically necessitates fields that couple strongly to the Higgs sector. Furthermore, the axion-Higgs mixing is also constrained by experimental bounds on rare meson decays Adler et al. (2004); Artamonov et al. (2009); Cortina Gil et al. (2021); Wei et al. (2009); Aaij et al. (2013, 2015) and beam-dump searches Bergsma et al. (1985) for axion-like particles. A comprehensive summary of these constraints can be found in Refs. Choi and Im (2016); Flacke et al. (2017). The bound from NA62 Cortina Gil et al. (2021) has been incorporated, updating the previous analyses Jeong et al. (2020, 2019). These combined constraints impose
@@
if the axion mass lies between approximately $300$ MeV and $5$ GeV. Here ${\rm Br}(\phi\to\mu^{+}\mu^{-})$ denotes the branching ratio for the axion decay into a muon pair. Additionally, if the axion mass is below the supernova core temperature, the mixing should be sufficiently suppressed to avoid accelerating supernova cooling. Fig. 1 provides a summary of the constraints on axion-Higgs mixing under the assumption that the axion decays exclusively into SM particles. The light brown band illustrates the relation, $\sin^{2}\theta_{\rm mix}=\kappa_{\rm mix}(m_{\phi}/m_{h})^{2}$, where $\kappa_{\rm mix}=(\kappa_{2}/\kappa_{1})^{2}$ is expressed in terms of the model parameters as
@@
significant constraints arise from supernovae explosions, neutron star merger events, and the predictions of standard BBN, particularly for the values of $f_{\phi}$ required for the successful baryogenesis. The SN 1987A neutrino signal imposes the constraint
@@
unless $m_{\phi}\gtrsim 200$ MeV Caputo et al. (2022a). Other constraints arise if the axion predominantly decay into SM particles, and these can be summarized as follows. Let us summarize such constraints. First, if axions escape freely from the SN core, the SN 1987A $\gamma$-ray signal Caputo et al. (2022a, b); Hoof and Schulz (2023); Müller et al. (2023) and the diffuse cosmic $\gamma$-ray background place stringent limits on the axion coupling to photons Caputo et al. (2022a, b). Additionally, axions produced in the protoneutron star could decay into photons, forming a fireball, which is a plasma of charged particles and photons Diamond et al. (2023). The absence of a fireball signal in the energy range $0.2$ and $2\ {\rm MeV}$ from the Pioneer Venus Orbiter (PVO) satellite provides further constraints on the axion-photon coupling. Second, axions should not deposit more energy in the SN mantle than is consistent with observations Caputo et al. (2022b). Furthermore, a recent study has revealed that axions can be copiously produced in neutron star mergers, with their decay leading to fireball formation Diamond et al. (2024). The non-observation of such a fireball following the emission of GW170817 Abbott et al. (2017) places additional constraints on the axion-photon coupling. If the axion lifetime is sufficiently long to prevent fireball formation, the resulting $\gamma$-rays would still be constrained by Fermi-LAT observations Dev et al. (2024). We note that the aforementioned astrophysical constraints that arise when the axion dominantly decays to SM particles can be alleviated in the presence of light dark particles that couple to the axion. This possibility will be discussed in detail in subsequent sections. The BBN constraints apply for values of $g_{\phi\gamma}$ smaller than about $3\times 10^{-9}\,{\rm GeV}^{-1}$ when the axion is heavier than $1$ MeV Cadamuro and Redondo (2012). However, the disruption of BBN can be evaded if the axion lifetime is sufficiently short, which is achievable through axion-Higgs mixing and the presence of light dark particles coupled to the axion. Lastly, for $g_{\phi\gamma}$ larger than $2\times 10^{-7}\,{\rm GeV}^{-1}$, beam dump experiments impose limits, depending on the axion mass O’Hare (2020).
@@ III Dark photons
Light dark particles coupled to the axion can alleviate the constraints on the axion-photon coupling imposed by BBN and supernova explosions by enhancing the axion decay rate while simultaneously suppressing the branching ratio of axion decay into SM particles. Since axion interactions are governed by the global U$(1)_{\phi}$ symmetry, dark photons, whose coupling to the axion naturally arises from the U$(1)_{\phi}$ anomaly, emerge as a compelling candidate for these light dark particles.444 Although we do not consider this possibility in detail here, an interesting alternative involves the inclusion of dark fermions coupled to the axion. Their interactions are constrained by U$(1)_{\phi}$, and are generally of two types, given by $\frac{k_{1}\partial_{\mu}\phi}{f_{\phi}}\bar{f}\gamma^{\mu}\gamma_{5}f+m_{f}e^{ik_{2}\frac{\phi}{f_{\phi}}}\bar{f}f,$ where $m_{f}$ denotes the mass of $f$, and $k_{1}$ and $k_{2}$ are constants of order unity. Physical quantities depend on a combination of $k_{1}$ and $k_{2}$ that remains invariant under chiral field rotations. These interactions give rise to $y_{f}\phi\bar{f}f$, with a Yukawa coupling given by $y_{f}\sim m_{f}/f_{\phi}$, indicating that the axion couplings to dark fermions are more constrained compared to those with dark photons. Axion-coupled dark photons can also play an important role in cosmology Agrawal et al. (2018); Kitajima et al. (2018); Agrawal et al. (2020); Kitajima et al. (2022); Nakagawa et al. (2023), with their impact depending on their masses generated through either the Higgs or Stueckelberg mechanism Stueckelberg (1938). Although not considered in this work, dark photons may exhibit kinetic mixing with ordinary photons. This kinetic mixing, which should be sufficiently small to remain consistent with various constraints Fabbrichesi et al. (2020); Caputo et al. (2021), would allow dark photons to interact with SM particles.
@@ III.1 Invisible axion decay
The constraints from supernova explosions and the neutron merger for GW170817 are sensitive to the branching ratio of axion decay into SM particles. To suppress the branching ratio, the axion coupling to dark photons should be much larger than the coupling to electrons, leading to the condition
$g_{\phi\gamma^{\prime}_{\rm M}}\gg\frac{2\kappa_{2}m_{e}}{m_{\phi}f_{\phi}}\simeq\left(\frac{\kappa_{2}}{\kappa_{1}}\right)\times 10^{-7}\,{\rm GeV}^{-1},$ (III.7)
@@
relaxing the constraints from supernova explosions and the neutron merger for GW170817. The branching ratio can be further suppressed by introducing additional dark photons.
@@
$g_{\phi\gamma^{\prime}_{M}}>4\times 10^{-9}\,{\rm GeV}^{-1}\left(\frac{m_{\phi}}{10\,{\rm MeV}}\right)^{-1},$ (III.9)
@@
$\sin^{2}\theta_{\rm mix}>10^{-11}.$ (III.10)
In our model, the axion-Higgs mixing is already sufficiently large to meet this requirement, except in the case where $\kappa_{2}$ is extremely small. Dark photons with a coupling given by Eq. (III.7) further reduce the axion lifetime. This indicates that the primary role of dark photons is to alleviate the astrophysical constraints arising from supernova explosions.
@@
$\Gamma_{\rm ID}\approx\frac{n_{\phi}}{n_{\gamma^{\prime}_{\rm M}}}\frac{1}{64\pi}g^{2}_{\phi\gamma^{\prime}_{\rm M}}m^{3}_{\phi},$ (III.11)
falls below the Hubble expansion rate. Here, $n_{i}$ is the number density of the species $i$, and we take the dark photon mass to be $23\,{\rm MeV}<m_{\gamma^{\prime}_{\rm M}}<m_{\phi}/$ 2, with the lower bound ensuring that the dark photon becomes non-relativistic prior to BBN. After freeze-out, dark photons dominantly decay into ultralight dark fermions. The contribution of these dark fermions to the effective number neutrino species can be estimated as
$\Delta N_{\rm eff}\approx\left(\frac{T_{\rm fo}}{T_{\rm decay}}\right)\left(\frac{m_{\gamma^{\prime}_{\rm M}}}{T_{\rm fo}}\right)^{5/2}e^{-m_{\gamma^{\prime}_{\rm M}}/T_{\rm fo}},$ (III.12)
under the assumption that $T_{\rm decay}<T_{\rm fo}\lesssim m_{\gamma^{\prime}_{\rm M}}/3$, where $T_{\rm fo}$ and $T_{\rm decay}$ denote the freeze-out and decay temperature of the dark photon, respectively. Although the out-of equilibrium decay of axions also produces dark fermions, their contribution is negligible due to the large Boltzmann suppression of the axion density compared to that of the dark photon. The freeze-out temperature $T_{\rm fo}$ is determined by the condition $\Gamma_{\rm ID}\approx H$ at $T=T_{\rm fo}$, which leads to
$\frac{m_{\phi}-m_{\gamma^{\prime}_{\rm M}}}{T_{\rm fo}}$$\approx$ $11+2\ln\left(\frac{m_{\phi}/T_{\rm fo}}{10}\right)$ (III.13)
@@
${\cal L}=\varepsilon A^{\prime}_{\mu}\bar{\psi}\gamma^{\mu}\psi,$ (III.14)
@@
$\varepsilon<10^{-9}\left(\frac{m_{\gamma^{\prime}_{\rm M}}}{10\,{\rm MeV}}\right)^{1/2},$ (III.15)
@@ III.2 Dark matter
$T_{\rm reh}<10^{6}\,{\rm GeV}\left(\frac{g_{\phi\gamma^{\prime}_{\rm DM}}}{10^{-9}\,{\rm GeV}^{-1}}\right)^{-4/3},$ (III.16)
@@
$g_{\phi\gamma^{\prime}_{\rm DM}}<10^{-8}\,{\rm GeV}^{-1}\left(\frac{m_{\phi}}{10\,{\rm MeV}}\right)^{-1/2},$ (III.17)
@@
$\Omega_{\gamma^{\prime}_{\rm DM}}$$\approx$ $\Omega_{\rm DM}\left(\frac{m_{\gamma^{\prime}_{\rm DM}}}{100\,{\rm keV}}\right)\left(\frac{g_{\phi\gamma^{\prime}_{\rm DM}}}{10^{-10}\,{\rm GeV}^{-1}}\right)^{2}$ (III.18)
@@
under the assumption that the initial dark photon abundance is negligibly small. Here we denote the observed density parameter for dark matter as $\Omega_{\rm DM}$, and $g_{\ast}(T)$ is taken to be the value at a temperature around $m_{\phi}$, the point at which the freeze-in process is most efficient. This expression indicates that massive dark photons can constitute the dominant component of dark matter in the universe across a wide range of the parameter space. We have confirmed that the approximate relation above is consistent with the dark photon relic density obtained from numerically solving the Boltzmann equation. Furthermore, we note that the condition (III.17), necessary to prevent the thermalization of dark photons, is always satisfied within the parameter space where the dark photon density does not exceed the observed dark matter density. In addition, the dark photon dark matter can be produced from the quantum fluctuation during inflation Graham et al. (2016); Ema et al. (2019). The abundance is sufficiently suppressed if the Hubble parameter during inflation is below approximately $10^{9}$ GeV.
@@ III.3 Dark radiation
An ultralight dark photon, denoted as $\gamma^{\prime}_{\rm DR}$, with mass $m_{\gamma^{\prime}_{\rm DR}}\lesssim 1$ eV and coupling to the axion represented by $g_{\phi\gamma^{\prime}_{\rm DR}}$, and behaves as dark radiation because dark photons are produced from thermal bath through the processes $\phi\phi\to\gamma^{\prime}_{\rm DR}\gamma^{\prime}_{\rm DR}$ and $\phi\to\gamma^{\prime}_{\rm DR}\gamma^{\prime}_{\rm DR}$. Dark photons get thermalized via the scattering process $\phi\phi\leftrightarrow\gamma^{\prime}_{\rm DR}\gamma^{\prime}_{\rm DR}$ if its decoupling temperature $T_{\rm dec\,{\rm DR}}$ is lower then $T_{\rm reh}$. It should be noted that the condition $T_{\rm dec\,{\rm DR}}<T_{\rm reh}$ is only applicable for dark photons with mass below the eV scale. Otherwise, freeze-out dark photons would result in overclosure of the universe. Upon decoupling from thermal bath, dark photons contribute to the effective number of neutrino species as follows
$\Delta N_{\rm eff}=\frac{4g}{7}\left(\frac{11}{4}\right)^{4/3}\left(\frac{g_{\ast s}(T)}{g_{\ast s}(T_{\rm dec\,DR})}\right)^{4/3},$ (III.19)
@@
$T_{\gamma^{\prime}_{\rm DR}}=\left(\frac{g_{\ast s}(T)}{g_{\ast s}(T_{\rm dec\,DR})}\right)^{1/3}T,$ (III.20)

### v3

@@ II.2 Constraints on axion
This contribution remains below the ACME II and JILA limits Andreev et al. (2018); Roussy et al. (2023) for $f_{\phi}$ larger than about $10$ TeV. Notably, many other electroweak baryogenesis scenarios face stringent EDM constraints, as achieving a strong first-order phase transition typically necessitates fields that couple strongly to the Higgs sector. Furthermore, the axion-Higgs mixing is also constrained by experimental bounds on rare meson decays Adler et al. (2004); Artamonov et al. (2009); Cortina Gil et al. (2021); Wei et al. (2009); Aaij et al. (2013, 2015) and beam-dump searches Bergsma et al. (1985) for axion-like particles. Additionally, if the axion mass is below the SN core temperature, axions can be produced abundantly in a core-collapse SN via interactions with nuclear matter. To prevent excessive SN cooling, the axion-Higgs mixing should either be sufficiently small to suppress axion production or sufficiently large to trap axions within the SN core Raffelt and Seckel (1988); Burrows et al. (1990); Carenza et al. (2019). In our model, since successful baryogenesis requires a decay constant below $10^{7}$ GeV, the axion should reside in the trapping regime. A comprehensive summary of these constraints can be found in Refs. Choi and Im (2016); Flacke et al. (2017); Winkler (2019). The bound from NA62 Cortina Gil et al. (2021) has been incorporated, updating the previous analyses Jeong et al. (2020, 2019). These combined constraints impose
@@
if the axion mass lies between approximately $300$ MeV and $5$ GeV. Here ${\rm Br}(\phi\to\mu^{+}\mu^{-})$ denotes the branching ratio for the axion decay into a muon pair. Fig. 1 provides a summary of the constraints on axion-Higgs mixing under the assumption that the axion decays exclusively into SM particles. The light brown band illustrates the relation, $\sin^{2}\theta_{\rm mix}=\kappa_{\rm mix}(m_{\phi}/m_{h})^{2}$, where $\kappa_{\rm mix}=(\kappa_{2}/\kappa_{1})^{2}$ is expressed in terms of the model parameters as
@@
significant constraints arise from SN explosions, neutron star merger events, and the predictions of standard BBN, particularly for the values of $f_{\phi}$ required for the successful baryogenesis. The SN 1987A neutrino signal imposes the constraint
@@
for $m_{\phi}\lesssim 200$ MeV Caputo et al. (2022a). However, this bound can be avoided if axions strongly interact with nuclear matter via axion-Higgs mixing, effectively trapping them in the SN core. Even if the energy loss from the SN via axion emission falls below the bound, additional constraints arise when the axion predominantly decay into SM particles. Let us summarize such constraints. First, if axions escape freely from the SN core, the SN 1987A $\gamma$-ray signal Caputo et al. (2022a, b); Hoof and Schulz (2023); Müller et al. (2023) and the diffuse cosmic $\gamma$-ray background place stringent limits on the axion coupling to photons Caputo et al. (2022a, b). Additionally, axions produced in the protoneutron star could decay into photons, forming a fireball, which is a plasma of charged particles and photons Diamond et al. (2023). The absence of a fireball signal in the energy range $0.2$ and $2\ {\rm MeV}$ from the Pioneer Venus Orbiter (PVO) satellite provides further constraints on the axion-photon coupling. Second, axions should not deposit more energy in the SN mantle than is consistent with observations Caputo et al. (2022b). Furthermore, a recent study has revealed that axions can be copiously produced in neutron star mergers, with their decay leading to fireball formation Diamond et al. (2024). The non-observation of such a fireball following the emission of GW170817 Abbott et al. (2017) places additional constraints on the axion-photon coupling. If the axion lifetime is sufficiently long to prevent fireball formation, the resulting $\gamma$-rays would still be constrained by Fermi-LAT observations Dev et al. (2024).
We note that the aforementioned astrophysical constraints that arise when the axion dominantly decays to SM particles can be alleviated in the presence of light dark particles that couple sufficiently strongly to the axion, while remaining weak enough to avoid disrupting the trapping of axions within the SN core. This possibility will be discussed in detail in subsequent sections. The BBN constraints apply for values of $g_{\phi\gamma}$ smaller than about $3\times 10^{-9}\,{\rm GeV}^{-1}$ when the axion is heavier than $1$ MeV Cadamuro and Redondo (2012). However, the disruption of BBN can be evaded if the axion lifetime is sufficiently short, which is achievable through axion-Higgs mixing and the presence of light dark particles coupled to the axion. Lastly, for $g_{\phi\gamma}$ larger than $2\times 10^{-7}\,{\rm GeV}^{-1}$, beam dump experiments impose limits, depending on the axion mass O’Hare (2020).
@@ III Dark photons
Light dark particles coupled to the axion can alleviate the constraints on the axion-photon coupling imposed by BBN and SN explosions by enhancing the axion decay rate while simultaneously suppressing the branching ratio of axion decay into SM particles. Since axion interactions are governed by the global U$(1)_{\phi}$ symmetry, dark photons, whose coupling to the axion naturally arises from the U$(1)_{\phi}$ anomaly, emerge as a compelling candidate for these light dark particles.444 Although we do not consider this possibility in detail here, an interesting alternative involves the inclusion of dark fermions coupled to the axion. Their interactions are constrained by U$(1)_{\phi}$, and are generally of two types, given by $\frac{k_{1}\partial_{\mu}\phi}{f_{\phi}}\bar{f}\gamma^{\mu}\gamma_{5}f+m_{f}e^{ik_{2}\frac{\phi}{f_{\phi}}}\bar{f}f,$ where $m_{f}$ denotes the mass of $f$, and $k_{1}$ and $k_{2}$ are constants of order unity. Physical quantities depend on a combination of $k_{1}$ and $k_{2}$ that remains invariant under chiral field rotations. These interactions give rise to $y_{f}\phi\bar{f}f$, with a Yukawa coupling given by $y_{f}\sim m_{f}/f_{\phi}$, indicating that the axion couplings to dark fermions are more constrained compared to those with dark photons. Axion-coupled dark photons can also play an important role in cosmology Agrawal et al. (2018); Kitajima et al. (2018); Agrawal et al. (2020); Kitajima et al. (2022); Nakagawa et al. (2023), with their impact depending on their masses generated through either the Higgs or Stueckelberg mechanism Stueckelberg (1938). Although not considered in this work, dark photons may exhibit kinetic mixing with ordinary photons. This kinetic mixing, which should be sufficiently small to remain consistent with various constraints Fabbrichesi et al. (2020); Caputo et al. (2021), would allow dark photons to interact with SM particles.
@@ III.1 Invisible axion decay
The constraints on the axion-photon coupling from SN explosions and the neutron merger for GW170817 are sensitive to the branching ratio of axion decay into SM particles. To suppress the branching ratio, the axion coupling to dark photons should be much larger than the coupling to electrons, leading to the condition
$g_{\phi\gamma^{\prime}_{\rm M}}>\frac{2\kappa_{2}m_{e}}{m_{\phi}f_{\phi}}\simeq\left(\frac{\kappa_{2}}{\kappa_{1}}\right)\times 10^{-7}\,{\rm GeV}^{-1},$ (III.7)
@@
relaxing the constraints from SN explosions and the neutron merger for GW170817. The branching ratio can be further suppressed by introducing additional dark photons. It is important to note that the axion coupling to a dark sector is constrained by the requirement that axions remain trapped within the SN core through their interactions with nuclear matter. This requires that the axion decay length should be much longer than the axion mean free path in the SN core, ensuring that axions are absorbed through the inverse pion conversion and inverse bremsstrahlung before they decay to dark particles Turner (1988); Caputo and Raffelt (2024). For the coupling constant given by
$g_{\phi\gamma^{\prime}_{\rm M}}<10^{-6}\,{\rm GeV}^{-1}\left(\frac{m_{\phi}}{10\,{\rm MeV}}\right)^{-3/2},$ (III.9)
the axion decay length at rest is larger than $\mathcal{O(}10)$ km, the size of the core, thereby not interfering with the trapping processes. We assume the existence of dark photons that couple to the axion, with coupling strengths constrained by (III.7) and (III.9).
@@
$g_{\phi\gamma^{\prime}_{M}}>4\times 10^{-9}\,{\rm GeV}^{-1}\left(\frac{m_{\phi}}{10\,{\rm MeV}}\right)^{-1},$ (III.10)
@@
$\sin^{2}\theta_{\rm mix}>10^{-11}.$ (III.11)
In our model, the axion-Higgs mixing is already sufficiently large to meet this requirement, except in the case where $\kappa_{2}$ is extremely small. Dark photons with a coupling given by Eq. (III.7) further reduce the axion lifetime. This indicates that the primary role of dark photons is to alleviate the astrophysical constraints arising from SN explosions.
@@
$\Gamma_{\rm ID}\approx\frac{n_{\phi}}{n_{\gamma^{\prime}_{\rm M}}}\frac{1}{64\pi}g^{2}_{\phi\gamma^{\prime}_{\rm M}}m^{3}_{\phi},$ (III.12)
falls below the Hubble expansion rate. Here, $n_{i}$ is the number density of the species $i$, and we take the dark photon mass to be $23\,{\rm MeV}<m_{\gamma^{\prime}_{\rm M}}<m_{\phi}/$ 2, with the lower bound ensuring that the dark photon becomes non-relativistic prior to BBN. After freeze-out, dark photons dominantly decay into ultralight dark fermions. The contribution of these dark fermions to the effective number of neutrino species can be estimated as
$\Delta N_{\rm eff}\approx\left(\frac{T_{\rm fo}}{T_{\rm decay}}\right)\left(\frac{m_{\gamma^{\prime}_{\rm M}}}{T_{\rm fo}}\right)^{5/2}e^{-m_{\gamma^{\prime}_{\rm M}}/T_{\rm fo}},$ (III.13)
under the assumption that $T_{\rm decay}<T_{\rm fo}\lesssim m_{\gamma^{\prime}_{\rm M}}/3$, where $T_{\rm fo}$ and $T_{\rm decay}$ denote the freeze-out and decay temperature of the dark photon, respectively. Although the out-of-equilibrium decay of axions also produces dark fermions, their contribution is negligible due to the large Boltzmann suppression of the axion density compared to that of the dark photon. The freeze-out temperature $T_{\rm fo}$ is determined by the condition $\Gamma_{\rm ID}\approx H$ at $T=T_{\rm fo}$, which leads to
$\frac{m_{\phi}-m_{\gamma^{\prime}_{\rm M}}}{T_{\rm fo}}$$\approx$ $11+2\ln\left(\frac{m_{\phi}/T_{\rm fo}}{10}\right)$ (III.14)
@@
${\cal L}=\varepsilon A^{\prime}_{\mu}\bar{\psi}\gamma^{\mu}\psi,$ (III.15)
@@
$\varepsilon<10^{-9}\left(\frac{m_{\gamma^{\prime}_{\rm M}}}{10\,{\rm MeV}}\right)^{1/2},$ (III.16)
@@ III.2 Dark matter
$T_{\rm reh}<10^{6}\,{\rm GeV}\left(\frac{g_{\phi\gamma^{\prime}_{\rm DM}}}{10^{-9}\,{\rm GeV}^{-1}}\right)^{-4/3},$ (III.17)
@@
$g_{\phi\gamma^{\prime}_{\rm DM}}<10^{-8}\,{\rm GeV}^{-1}\left(\frac{m_{\phi}}{10\,{\rm MeV}}\right)^{-1/2},$ (III.18)
@@
$\Omega_{\gamma^{\prime}_{\rm DM}}$$\approx$ $\Omega_{\rm DM}\left(\frac{m_{\gamma^{\prime}_{\rm DM}}}{100\,{\rm keV}}\right)\left(\frac{g_{\phi\gamma^{\prime}_{\rm DM}}}{10^{-10}\,{\rm GeV}^{-1}}\right)^{2}$ (III.19)
@@
under the assumption that the initial dark photon abundance is negligibly small. Here we denote the observed density parameter for dark matter as $\Omega_{\rm DM}$, and $g_{\ast}(T)$ is taken to be the value at a temperature around $m_{\phi}$, the point at which the freeze-in process is most efficient. This expression indicates that massive dark photons can constitute the dominant component of dark matter in the universe across a wide range of the parameter space. We have confirmed that the approximate relation above is consistent with the dark photon relic density obtained from numerically solving the Boltzmann equation. Furthermore, we note that the condition (III.18), necessary to prevent the thermalization of dark photons, is always satisfied within the parameter space where the dark photon density does not exceed the observed dark matter density. In addition, the dark photon dark matter can be produced from the quantum fluctuation during inflation Graham et al. (2016); Ema et al. (2019). The abundance is sufficiently suppressed if the Hubble parameter during inflation is below approximately $10^{9}$ GeV.
@@ III.3 Dark radiation
An ultralight dark photon, denoted as $\gamma^{\prime}_{\rm DR}$, with mass $m_{\gamma^{\prime}_{\rm DR}}\lesssim 1$ eV and coupling to the axion represented by $g_{\phi\gamma^{\prime}_{\rm DR}}$, and behaves as dark radiation because dark photons are produced from thermal bath through the processes $\phi\phi\to\gamma^{\prime}_{\rm DR}\gamma^{\prime}_{\rm DR}$ and $\phi\to\gamma^{\prime}_{\rm DR}\gamma^{\prime}_{\rm DR}$. Dark photons get thermalized via the scattering process $\phi\phi\leftrightarrow\gamma^{\prime}_{\rm DR}\gamma^{\prime}_{\rm DR}$ if its decoupling temperature $T_{\rm dec\,{\rm DR}}$ is lower than $T_{\rm reh}$. It should be noted that the condition $T_{\rm dec\,{\rm DR}}<T_{\rm reh}$ is only applicable for dark photons with mass below the eV scale. Otherwise, freeze-out dark photons would result in overclosure of the universe. Upon decoupling from thermal bath, dark photons contribute to the effective number of neutrino species as follows
$\Delta N_{\rm eff}=\frac{4g}{7}\left(\frac{11}{4}\right)^{4/3}\left(\frac{g_{\ast s}(T)}{g_{\ast s}(T_{\rm dec\,DR})}\right)^{4/3},$ (III.20)
@@
$T_{\gamma^{\prime}_{\rm DR}}=\left(\frac{g_{\ast s}(T)}{g_{\ast s}(T_{\rm dec\,DR})}\right)^{1/3}T,$ (III.21)